OpenStudy (anonymous):

Use the logarithmic properties to expand ln(x^2-16 divided by x^4), where x>4.

OpenStudy (anonymous):

\[(\frac{ x^2-16 }{ x^4 })\]

OpenStudy (anonymous):

\[\ln(\frac{ a }{ b })=lna-lnb\]

OpenStudy (anonymous):

\[\ln(a^n)=nlna\]

OpenStudy (anonymous):

ln(x^2-16)-ln(x^4)=ln(x^2-4^2)-ln(x^4)=ln[(x+4).(x-4)]-4.ln(x)=ln(x+4)+ln(x-4)-4.ln(x)

OpenStudy (inkyvoyd):

ok this is not ok. Someone else (probably you) just had the EXACT same question.

OpenStudy (inkyvoyd):

cheating on the same test I bet. Oh well.

OpenStudy (anonymous):

Is that Kelly? I didn't know she has a openstudy.